Tobac Co. is a monopolist in the cigarette market in Nicotiana Republic, where the U.S. dollar is used as the official currency. The firm faces the demand curve shown. The firm has a constant marginal cost of $2.00 per pack. The fixed cost of the firm is $50 million.

To answer the question, it is useful to know that the equation of the (inverse) demand curve is
P = 8 - 0.04Q, where Q is the quantity demanded (in millions of packs) and P is the price per pack (in $). Also, you could draw in the marginal revenue curve.
If the quantity sold is 150 million packs, the firm's profit is?
Answer #1

Hello,

Please find the answer below:

Price per pack = 8 - 0.04Q

TR = P*Q

Q = 150 million packs

TR = 8*150 - 0.04*(150)^2 = $ 300 million

MC = $ 2 per pack = $2 *150 million = $300 million

TC = 300 + 50 = $ 350 million

Profit = TR-TC

= 300 - 350 = $ (50) million

i.e. Loss of $ 50 million